### v1

# Self-referential instances of the dominating set problem are irreducible

Abstract
We study the algorithmic decidability of the domination number in the Erdős–Rényi random graph model $G(n,p)$. We show that for a carefully chosen edge probability $p=p(n)$, the domination problem exhibits a strong irreducible property. Specifically, for any constant $0<c<1$, no algorithm that inspects only an induced subgraph of order at most $n^{c}$ can determine whether $G(n,p)$ contains a dominating set of size $k=\ln n$. We demonstrate that the existence of such a dominating set can be flipped by a local symmetry mapping that alters only a constant number of edges, thereby producing indistinguishable random graph instances which requires exhaustive search. These results demonstrate that the extreme hardness of the dominating set problem in random graphs cannot be attributed to local structure, but instead arises from the self-referential nature and near-independence structure of the entire solution space.
1 Introduction
The domination set problem is a classic NP-hard problem in graph theory. Given an Erdős-Rényi random graph $G(n,p)$, that is, a graph on $n$ vertices in which each of the $\binom{n}{2}$ potential edges appears independently with probability $p$. A dominating set is a subset of vertices such that every vertex of $G(n,p)$ either belongs to the set or is adjacent to at least one vertex in the set. Despite extensive study, no algorithm substantially better than the trivial brute-force search, which checks all subsets of size $k$, is known in general. This difficulty has motivated research into alternative approaches, including fixed-parameter tractable (FPT) algorithms and approximation schemes [2, 1].
From the perspective of parameterized complexity, however, dominating set is known to be W[2]-hard, which strongly suggests that no algorithm with running time $f(k)n^{O(1)}$ exists unless the $W$- hierarchy collapses. In parallel, significant progress has been made in understanding the domination number of the random graphs. In particular, Wieland and Godbole [5] showed that for certain ranges of $p$, mirroring classical concentration phenomena observed for other graph parameters such as the independence number and the chromatic number. In contrast, when $p$ is small, for example $p=O(1/n)$, the domination number exhibits markedly different behavior: it is not concentrated on any interval of length $o(\sqrt{n})$, and the threshold for domination shifts from the first-moment bound to the median, with a substantial gap between the two [3].
In this paper, we establish a strong indistinguishability phenomenon for the domination set problem in random graphs. Specifically, for any constant $0<c<1$, we show that it is impossible to determine whether $G(n,p)$ contains a dominating set of size $k=\ln n$ by inspecting any induced subgraph $H$ of order at most $n^{c}$. To prove this result, we carefully choose the edge probability $p$ and show that there exist two types of instances of $G(n,p)$: one class consists of graphs with a unique dominating set of size $k$, while the other consists of graphs that contain no dominating set of size $k$, but instead contain vertex sets of size $k$ that dominate all but a single vertex.
A crucial feature of our construction is that the existence or non-existence of a dominating set of size $k$ can be toggled by altering the adjacency between four vertices. This change alters the existence of a dominating set of size $k$ while leaving the subgraph $H$ unchanged. Consequently, any algorithm whose decision is based solely on the information contained in $H$ cannot reliably distinguish between these two types of instances, and therefore cannot determine the existence of a dominating set of size $k$ without effectively examining the entire graph.
Our methodology is closely related to the symmetry-mapping technique introduced in the construction of self-referential (most indistinguishable) instances for Model RB in [6]. These instances were shown to require essentially brute-force computation for their solution. The present work extends this line of reasoning to the Dominating Set problem, providing further evidence that partial information cannot substitute for global structure. In particular, our results support the view that the fundamental source of extreme hardness in such problems lies in the self-referential nature and near-independence structure of their candidate solutions.
2 Main results
Definition 2.1.
Let $G=G(n,p)$ be a random graph. We say that the dominating set problem for $G$ is reducible if there exists an induced subgraph $H$ of order $n^{c}$, for some constant $0<c<1$, such that $H$ contains a dominating set of size $k$ if and only if $G$ contains a dominating set of size $k$; otherwise $G$ is said to be irreducible.
The following theorem shows that, under this definition, the dominating set problem in random graphs cannot be reduced to any subgraph of order $n^{c}$ for any constant $0<c<1$.
Theorem 2.1.
The dominating set problem for $G(n,p)$ is irreducible.
2.1 Existence of a unique dominating set of size $k$
In this section we show that, with positive probability, the random graph $G(n,p)$ contains a unique dominating sets of size $k=\ln n$. Let $X$ be the number of dominating sets of size $k$ in $G(n,p)$. A straightforward calculation yields
$\mathbf{E}[X]$ $=\binom{n}{k}(1-(1-p)^{k})^{n-k},$ (2.1)
$\mathbf{E}[X^{2}]$ $=\sum_{i=0}^{k}\binom{n}{i}\binom{n-i}{k-i}\binom{n-k}{k-i}(1-(1-p)^{k})^{2(k-i)}(1-2(1-p)^{k}+(1-p)^{2k-i})^{n-2k+i}.$ (2.2)
In the following, we tacitly choose the edge probability $p=p(n)$ so that $E[X]=\delta+o(1)$ for a constant $0<\delta<1$. Standard asymptotic analysis shows that this choice is achieved, for example, by
$p=1-\frac{1}{e}\bigl((1-\varepsilon_{n})\ln^{2}n\bigr)^{1/\ln n},\qquad\varepsilon_{n}=(1+o(1))\frac{\ln\ln n}{\ln n}.$ (2.3)
Lemma 2.2.
In $G(n,p)$,
$\frac{\delta}{\delta+1}\leq\mathbf{Pr}(X>0)\leq\delta.$
Proof.
The upper bound follows immediately from Markov’s inequality:
$\mathbf{Pr}(X>0)\leq\mathbf{E}[X]=\delta+o(1).$ (2.4)
For a lower bound we apply the second moment method. $\mathbf{E}[X^{2}]$ counts ordered pairs of dominating sets of size $k$. Let $F(i)$ denote the contribution from pairs whose intersection has size $i$, then
$\mathbf{E}[X^{2}]$ $=\sum_{i=0}^{k}F(i),$
where
$F(i)=\binom{n}{i}\binom{n-i}{k-i}\binom{n-k}{k-i}\big(1-(1-p)^{k}\big)^{2(k-i)}\left(1-2(1-p)^{k}+(1-p)^{2k-i}\right)^{n-2k+i}.$
If $i=0$, we have
$F(0)=\binom{n}{k}\binom{n-k}{k}\big(1-(1-p)^{k}\big)^{2n-2k}=(1+o(1))\mathbf{E}[X]^{2}.$
If $i=k$, then
$F(k)=\mathbf{E}[X].$
For $1\leq i\leq k-1$,
$\frac{F(i)}{\mathbf{E}[X]}=\frac{(n-k)!k!}{(n-2k+i)!(k-i)!^{2}i!}\frac{(1-(1-p)^{k})^{2(k-i)}(1-2(1-p)^{k}+(1-p)^{2k-i})^{n-2k+i}}{(1-(1-p)^{k})^{n-k}}.$
[…]
$\frac{F(i)}{\mathbf{E}[X]}$ $\leq(1+o(1))n^{k-i}\frac{\bigl(1-2(1-p)^{k}+o((1-p)^{k})\bigr)^{n-2k+i}}{\bigl(1-(1-p)^{k}\bigr)^{n-k}}$
$=(1+o(1))\exp\bigl\{-2\ln^{2}n+(k-i)\ln n\bigr\}.$
Summing over $i$ and dividing by $\mathbf{E}[X]^{2}$, we obtain
$\frac{\mathbf{E}[X^{2}]}{\mathbf{E}[X]^{2}}\leq 1+\frac{1}{\mathbf{E}[X]}+o(1)=1+\frac{1}{\delta}+o(1).$ (2.5)
Finally, by the Cauchy-Schwarz inequality,
$\mathbf{Pr}(X>0)\geq\mathbf{E}[X]^{2}/\mathbf{E}{[X^{2}]}\geq\frac{\delta}{\delta+1}+o(1),$
which completes the proof. ∎
As an immediate consequence of Lemma 2.2 we obtain a lower bound of the probability that $G(n,p)$ has a unique dominating set of size $\ln n$.
Corollary 2.3.
The probability that $G(n,p)$ has a unique dominating set of size $\ln n$ is at least $\delta(1-\delta)/(1+\delta)$.
Proof.
Let $\rho_{1}$ be the probability that $G(n,p)$ has a unique dominating set of size $\ln n$, and $\rho_{\geq 2}$ be the probability that $G(n,p)$ has at least two such dominating sets. Then from Lemma 2.2, we have
$\ \mathbf{E}[X]=\delta\geq\rho_{1}+2\rho_{\geq 2},$
$\mathbf{Pr}(X>0)=\rho_{1}+\rho_{\geq 2}\geq\frac{\delta}{\delta+1}.$
Therefore $\rho_{1}\geq\delta(1-\delta)/(1+\delta)$. ∎
2.2 Non-existence of dominating sets of size $k$
Lemma 2.4.
In the random graph $G(n,p)$, if there does not exist a dominating set of size $k=\ln n$, then with high probability there exist $k$-vertex sets that dominates all but exactly one vertex.
Proof.
Let $S$ be a $k$-vertex set. We say $S$ dominates all but one vertex if exactly one vertex in $V\setminus S$ has no neighbor in $S$, while the remaining nk1n-k-1 vertices in $V\setminus S$ are adjacent to at least one vertex of $S$.
Let $N=\sum_{|S|=k}I_{S}$, where $I_{S}=\mathbf{1}_{\{\text{$S$ dominates all but one vertex}\}}$. A direct calculation gives
$\mathbf{E}[N]=\binom{n}{k}(n-k)(1-p)^{k}\big(1-(1-p)^{k}\big)^{n-k-1}.$ (2.6)
We compare $\mathbf{E}[N]$ with $\mathbf{E}[X]$ (2.1). Since
$\frac{\mathbf{E}[N]}{\mathbf{E}[X]}=(n-k)(1-p)^{k}\bigl(1-(1-p)^{k}\bigr)^{-1},$
and under our choice
$p=1-\frac{1}{e}\bigl((1-\varepsilon_{n})\ln^{2}n\bigr)^{1/\ln n},\qquad(1-p)^{k}=\frac{\ln^{2}n}{n}(1+o(1)),$
we obtain
$\mathbf{E}[N]=(1+o(1))(\ln n)^{2}\mathbf{E}[X]\to\infty.$
We now apply the second moment method. Writing
$\mathbf{E}[N^{2}]=\sum_{|S|=k}\sum_{|S^{\prime}|=k}\mathbf{E}[I_{S}I_{S^{\prime}}].$
Let $i=|S\cap S^{\prime}|$, then
$\mathbf{E}[N^{2}]=\sum_{i=0}^{k}\Phi(i)W(i),$
where
$\Phi(i)=\binom{n}{i}\binom{n-i}{k-i}\binom{n-k}{k-i},W(i)=\mathbf{E}\left[I_{S}I_{S^{\prime}}\bigm||S\cap S^{\prime}|=i\right]=P_{1}(i)+2P_{2}(i)+P_{3}(i)+P_{4}(i),$
and (the derivation of $W(i)$ is in the Appendix)
$P_{1}(i)$ $=(k-i)^{2}\mu^{2}(1-\mu)^{2k-2i-2}\zeta^{m},$
$P_{2}(i)$ $=(k-i)m\mu^{2}\vartheta(1-\mu)^{2k-2i-1}\zeta^{m-1},$
$P_{3}(i)$ $=m\varpi(1-\mu)^{2(k-i)}\zeta^{m-1},$
$P_{4}(i)$ $=m(m-1)(\mu\vartheta)^{2}(1-\mu)^{2k-2i}\zeta^{m-2},$
with $\mu\triangleq(1-p)^{k},\vartheta\triangleq 1-(1-p)^{k-i},\varpi\triangleq(1-p)^{2k-i},m\triangleq n-2k+i,\zeta\triangleq 1-2u+\varpi.$
Now we can rewrite 2.6 as
$\mathbf{E}[N]=(n-k)\binom{n}{k}\mu(1-\mu)^{n-k-1}.$
We estimate $\mathbf{E}[N^{2}]/\mathbf{E}[N]^{2}$ by considering three cases.
Case 1: $i=0$. Here
$\Phi(0)$ $=\binom{n}{k}\binom{n-k}{k}$,
$W(0)$ $=\big(k^{2}+2k(n-2k)(1-\mu)^{-1}+(n-2k)+(n-2k)(n-2k-1)\big)\mu^{2}(1-\mu)^{2n-2k-2}.$
A direct computation yields
$\frac{W(0)}{(n-k)^{2}\mu^{2}(1-\mu)^{2n-2k-2}}$ $=\frac{k^{2}+2k(n-2k)(1-\mu)^{-1}+(n-2k)+(n-2k)(n-2k-1)}{(n-k)^{2}}$
$=1+o(1),$
hence
$\frac{\Phi(0)W(0)}{\mathbf{E}[N]^{2}}=1+o(1).$
Case 2: $i=k$. In this case $S=S^{\prime}$, so
$\Phi(k)=\binom{n}{k},\quad W(k)=(n-k)\mu(1-\mu)^{n-k-1},$
thus $\Phi(k)W(k)=\mathbf{E}[N]$. Therefore,
$\frac{\Phi(k)W(k)}{\mathbf{E}[N]^{2}}=\frac{1}{\mathbf{E}[N]}=o(1).$
Case 3: $1\leq i\leq k-1$. It is easy to see that
$\frac{\Phi(i)}{\binom{n}{k}^{2}}\leq\frac{k^{2i}}{n^{i}}.$
Moreover, note that $k=\ln n$, standard asymptotic estimates give
$\mu=(1+o(1))\frac{\ln^{2}n}{n},\qquad\vartheta=1-(1-p)^{k-i}=1-o(1),$
and
$\zeta^{m}=(1-2(1-p)^{k}+(1-p)^{2k-i})^{m}\leq(1-\mu)^{2m}\exp\{(1+o(1))\ln^{2}n\}.$
Therefore,
$\frac{P_{1}(i)}{(n-k)^{2}\mu^{2}(1-\mu)^{2n-2k-2}}$ $=\frac{(k-i)^{2}}{(n-k)^{2}}(1-\mu)^{-2m}\zeta^{m}=o(1).$
$\frac{P_{2}(i)}{(n-k)^{2}\mu^{2}(1-\mu)^{2n-2k-2}}$ $=\frac{2(k-i)m}{(n-k)^{2}}\vartheta(1-\mu)^{-2m-1}\zeta^{m-1}=o(1).$
$\frac{P_{3}(i)}{(n-k)^{2}\mu^{2}(1-\mu)^{2n-2k-2}}$ $=\frac{m}{(n-k)^{2}}(1-p)^{-i}(1-\mu)^{-2m}\zeta^{m-1}=o(1).$
$\frac{P_{4}(i)}{(n-k)^{2}\mu^{2}(1-\mu)^{2n-2k-2}}$ $=\frac{m(m-1)}{(n-k)^{2}}\vartheta^{2}(1-\mu)^{-2m}\zeta^{m-2}=\exp\{(1+o(1))\ln^{2}n\}.$
A detailed calculation shows that
$\frac{\Phi(i)W(i)}{\mathbf{E}^{2}[N]}$ $=\frac{\Phi(i)}{\binom{n}{k}^{2}}\frac{P_{1}(i)+2P_{2}(i)+P_{3}(i)+P_{4}(i)}{(n-k)^{2}\mu^{2}(1-\mu)^{2n-2k-2}}$
$=(1+o(1))\frac{k^{2i}}{n^{i}}\exp\{(1+o(1))\ln^{2}n\}.$
Summing over all three cases, we obtain
$\frac{\mathbf{E}[N^{2}]}{\mathbf{E}^{2}[N]}\leq 1+o(1).$
The second moment method yields
$\mathbf{P}[N>0]\geq\frac{\mathbf{E}^{2}[N]}{\mathbf{E}[N^{2}]}\geq 1-o(1).$
Thus with high probability there exists at least one set of $k$ vertices that dominates all but exactly one vertex, completing the proof. ∎
3 Proof of Theorem 2.1
Let $\mathscr{G}$ denote the family of instances of $G(n,p)$ such that each instance either has a unique dominating set of size $k=\ln n$ or has no dominating set of that size. By Lemma 2.2 and Lemma 2.4, the probability that a random graph $G(n,p)$ belongs to $\mathscr{G}$ is bounded away from zero.
We show that for any constant $0<c<1$, no subgraph of order at most $n^{c}$ suffices to decide whether $G$ contains a dominating set of size $k$.
Case 1: $G$ has a unique dominating set of size $k$.
Assume that $G\in\mathscr{G}$ contains a unique dominating set $S$ of size $k$. By Corollary 2.3, this occurs with probability at least $1/6$. Let $H$ be an arbitrary induced subgraph of $G$ on at most $n^{c}$ vertices, and let $V_{G},V_{H}$ be the vertices in $G,H$ respectively.
Then
$\mathbf{P}(S\subseteq V_{G}\setminus V_{H})=\frac{\binom{n-n^{c}}{k}}{\binom{n}{k}}=1-o(1),$
so with high probability the vertices of $S$ lie entirely outside $H$.
We first show that with high probability there exists a vertex in $V_{G}\setminus(V_{H}\cup S)$ adjacent to exactly one vertex of $S$. For a vertex $v\notin S$, let
[…]
Then
$\mathbf{P}(A_{v}\mid B_{v})=\frac{\mathbf{P}(A_{v})}{\mathbf{P}(B_{v})}=\frac{kp(1-p)^{k-1}}{1-(1-p)^{k}}.$
Consequently,
$\mathbf{P}\!\left(\exists\,v\in V_{G}\setminus(V_{H}\cup S):A_{v}\right)=1-\left(1-\frac{kp(1-p)^{k-1}}{1-(1-p)^{k}}\right)^{n-n^{c}-k}=1-o(1).$
Fix such a vertex $v$, and let $u\in S$ be its unique neighbor in $S$. Next, choose vertices $z\in S$ and $w\in V_{G}\setminus(V_{H}\cup S)$ such that, among these vertices, the only existing edges are $(u,v)$ and $(z,w)$.
We now perform a symmetry mapping by removing the edges $(u,v)$ and $(z,w)$, and adding the edges $(u,z)$ and $(v,w)$, as illustrated in Figure 1 (transition from (a) to (b)). This operation modifies only edges incident to vertices outside $H$, while preserving the degree of every vertex and keeping the total number of edges unchanged. As a result of this transformation, vertex $v$ no longer has any neighbor in $S$, implying that $S$ is no longer a dominating set.
Moreover, with high probability no new dominating set of size $k$ is created. Indeed, the probability that a set containing $\{v\}$ or $\{w\}$ extends to a dominating set of size $k$ is at most
$\binom{n-n^{c}}{k-1}\bigl(1-(1-p)^{k}\bigr)^{n-n^{c}-k-1}=o(1).$
Thus, with high probability, the modified graph contains no dominating set of size $k$.
Figure 1: A symmetry mapping between two classes of instances.
(a) → (b) Initially, the vertex $v$ has exactly one neighbor $u\in S$. The symmetry mapping transforms $G$ which has a unique dominating set of size $k$ into a graph with no dominating set of size $k$.
(b) → (a) Initially, $v$ is the only vertex not dominated by $S$. The symmetry mapping transforms $G$ which has no dominating set of size $k$ into a graph with a unique dominating set of size $k$.
Case 2: $G$ has no dominating set of size $k$.
Now suppose that $G\in\mathscr{G}$ has no dominating set of size $k$. By Lemma 2.4, with high probability there exists a set $S$ of $k$ vertices and a vertex $v$ such that $S$ dominates all vertices except $v$.
Let $H$ be any induced subgraph of order at most $n^{c}$. Then
$\mathbf{P}(S\cup\{v\}\subseteq G\setminus H)=\frac{\binom{n-n^{c}}{k+1}}{\binom{n}{k+1}}=1-o(1),$
so with high probability all vertices in $S\cup\{v\}$ lie outside $H$.
Choose vertices $u,z\in S$ and $w\in V_{G}\setminus(V_{H}\cup S)$ such that, among these vertices, the only existing edges are $(v,w)$ and $(u,z)$. Apply the symmetry mapping in the opposite direction by removing the edges $(v,w)$ and $(u,z)$ and adding the edges $(v,u)$ and $(z,w)$. After this change, the vertex $v$ becomes adjacent to $S$, and hence $S$ forms a dominating set of size $k$. See Figure 1(from (b) to (a)).
As before, with high probability no other dominating set of size $k$ is created, since
$\binom{n-n^{c}}{k-1}\bigl(1-(1-p)^{k}\bigr)^{n-n^{c}-k-1}=o(1).$
Therefore, with high probability, the modified graph contains a unique dominating set of size $k$.
In both cases, by altering only edges whose endpoints lie outside $H$, we can flip the existence of a dominating set of size $k$ while keeping the induced subgraph $H$ unchanged. Consequently, for any $0<c<1$, no subgraph of order $n^{c}$ contains sufficient information to determine whether $G(n,p)$ has a dominating set of size $k$.
This proves that the dominating set problem for $G(n,p)$ is irreducible, and completes the proof of Theorem 2.1.
4 Conclusion
Our results provide another example of self-referential instances that inherently require exhaustive search, in the sense that no proper substructure contains sufficient information to decide the global property. In such instances, the existence of solutions cannot be inferred from any sublinear portion of the input, reflecting a fundamental obstruction to local or reduction-based algorithms. We argue that this phenomenon is closely related to the presence of independent solution spaces in the underlying combinatorial problem, where candidate solutions interact only weakly and local modifications can drastically alter global feasibility.
This mechanism naturally arises in certain constraint satisfaction problems. In the dominating set problem studied here, for example, two randomly chosen candidate vertex sets of size $n^{c}$ are disjoint with high probability, implying that their domination properties are nearly independent. This independence enables the construction of symmetry mappings that preserve all local subgraphs while flipping the existence of global solutions, thereby producing most indistinguishable instances. Similar phenomena have been observed in other classical problems. Notably, for the CLIQUE problem, Li et al. [4] constructed self-referential instances that expose an intrinsic source of algorithmic hardness beyond local graph structure.
Taken together, these results suggest that by understanding how self-reference and solution-space independence forces exhaustive search in specific cases, one can gain insight into the nature of extreme hardness, making “proving computational hardness not hard in mathematics” for these particular examples. This aligns with Gödel’s use of self-reference in logic to prove the existence of formally unprovable mathematical statements.
Appendix A Derivation of the function $W(i)$
Let $S,S^{\prime}$ be two sets, each of size $k$, that each dominate all but one vertex in $G(n,p)$. Suppose that $|S\cap S^{\prime}|=i$. We derive an explicit expression for
$W(i)=\mathbf{P}\big(I_{S}=1,I_{S^{\prime}}=1\bigm||S\cap S^{\prime}|=i\big),$
where $I_{S}=1$ (respectively $I_{S^{\prime}}=1)$ denotes the event that $S$ (resp. $S^{\prime}$) dominates all but one vertex.
Define the disjoint vertex sets
$A=S\cap S^{\prime},B=S\backslash S^{\prime},C=S^{\prime}\backslash S,D=V\backslash(S\cup S^{\prime}).$
Then
$|A|=i,|B|=|C|=k-i,|D|=n-2k+i.$
The events $I_{S}=1$ and $I_{S^{\prime}}=1$ can equivalently restated as follows:
  • $I_{S}=1$: Among the vertices in $C\cup D$, exactly one vertex has no neighbor in $S$;
  • $I_{S^{\prime}}=1$: Among the vertices in $B\cup D$, exactly one has no neighbor in $S^{\prime}$.
Let $x\in C\cup D$ be the unique vertex not dominated by $S$, and $y\in B\cup D$ be the unique vertex not dominated by $S^{\prime}$. We consider all possible locations of the pair $(x,y)$.
For convenience, introduce the notation
$\mu\triangleq(1-p)^{k},\vartheta\triangleq 1-(1-p)^{k-i},\varpi\triangleq(1-2p)^{2k-i},m\triangleq n-2k+i,\zeta\triangleq 1-2\mu-\varpi.$
Case 1: $x\in C$ and $y\in B$
The vertex $x$ has no neighbor in $S$ with probability $\mu$, and $y$ has no neighbor in $S^{\prime}$ with probability $\mu$. Each of the remaining ki1k-i-1 vertices in $C$ must be dominated by $S$, and each of the remaining ki1k-i-1 vertices in $B$ must be dominated by $S^{\prime}$, each occurring with probability $1-\mu$. Each vertex in $D$ must be dominated by both $S$ and $S^{\prime}$, which happens with probability $\zeta$. Since there are kik-i choices for xxin $C$ and kik-i choices for $y$ in $B$, we obtain
$P_{1}(i)=(k-i)^{2}\mu^{2}(1-\mu)^{2k-2i-2}\zeta^{m}.$
Case 2: $x\in C,y\in D$ (the symmetric case $x\in D,y\in B$ yields the same contribution).
Here $x$ has no neighbor in $S$, which occurs with probability $\mu$. The vertex $y$ has no neighbor in $S^{\prime}$ but is dominated by $S$, so it has no edges to $A\cup C$ and at least one edge to $B$; this occurs with probability $\mu\vartheta$. The remaining ki1k-i-1 vertices in $C$ are dominated by $S$, and all kik-i vertices in $B$ are dominated by $S^{\prime}$, each with probability $1-\mu$. The remaining m1m-1 vertices in $D$ are dominated by both $S$ and $S^{\prime}$, each with probability $\zeta$. There are kik-i choices for $x$ and $m$ choices for $y$, giving
$P_{2}(i)=(k-i)m\mu^{2}\vartheta(1-\mu)^{2k-2i-1}\zeta^{m-1}.$
Case 3: $x\in D,y\in D$, and $x=y$
In this case the common vertex $x=y$ has no neighbors in either $S$ or $S^{\prime}$, which occurs with probability $\varpi$. The remaining m1m-1 vertices in $D$ are dominated by both $S$ and $S^{\prime}$, each with probability $q$. All kik-i vertices in $C$ are dominated by $S$, and all kik-i vertices in $B$ are dominated by $S^{\prime}$, each with probability $1-\mu$. Since there are $m$ choices for the common vertex, we obtain
$P_{3}(i)=m\varpi(1-\mu)^{2(k-i)}\zeta^{m-1}.$
Case 4: $x\in D,y\in D$ and $x\neq y$.
Here $x$ has no neighbor in $S$ but is dominated by $S^{\prime}$, which occurs with probability $\mu\vartheta$; symmetrically, $y$ has no neighbor in $S^{\prime}$ but is dominated by $S$, also with probability $\mu\vartheta$. The remaining m2m-2 vertices in $D$ are dominated by both $S$ and $S^{\prime}$, each with probability $\zeta$. All vertices in $C$ are dominated by $S$, and all vertices in $B$ are dominated by $S^{\prime}$, each with probability $1-\mu$. There are m(m1)m(m-1) ordered choices for $(x,y)$, yielding
$P_{4}(i)=m(m-1)(\mu\vartheta)^{2}(1-\mu)^{2k-2i}\zeta^{m-2}.$
Combining the above cases, we have
$W(i)=P_{1}(i)+2P_{2}(i)+P_{3}(i)+P_{4}(i).$